Thank you for helping me out.

Logarithm - EXERCISE:
There are 60 wolves in a certain location. Each year there are 8.5% more wolves than in the previous year. Use the population formula for continuous growth (P=P0,e^(rt)) to determine:
a)How many wolves are in the population after 6 3/4 years?
Answer to nearest whole number.
b)How long will it take the population to double
(to the closest whole number of years)?

Answer by edjones(8007)

P=P[0]*e^(rt)
.
a)
60*e^(.085*6.75)
=60*e^.57375
=60*1.7749
=106 wolves are in the population after 6 3/4 years.
.
b)
60*e^.085d=120
e^.085d=2
ln(e^.085d)=ln(2)
.085d=.6931
d=8 yrs
.
Ed